How many grams of nitrogen gas are in a 2 L container at 1.5 atm and 25 centigrade?


1
Expert's answer
2020-10-01T07:55:30-0400

Use Ideal Gas Law

"PV=nRT"

"n=\frac{PV}{RT}=\frac{1.5 atm\times 2L}{0.082 \frac{L\times atm}{mol\times K}\times (273+25)}=0.123 moles"


convert moles to grams

"0.123 moles (N_2)\times \frac{28.01 g (N_2)}{1 mole (N_2)}=3.4 g (N_2)"